FOREWORD

The present issue is the third volume of the Josai Mathematical Monograghs (JMM). JMM is founded by Graduate School of Science of Josai University (GSSJU), and appears yearly. JMM aims at formal as well as informal information on developments in mathematical research by the mathematical faculty of GSSJU and others. The Faculty of Science of Josai University has been publishing yearly the Science Bulletin of Josai University (SBJU). As a part of the preparation for the establishment of GSSJU, the Faculty of Science of Josai University published Special Issues of SBJU irregularly, from No. 1 (1995) to No. 5 (1997). Fot this reason, the publication of Special Issues of SBJU is the prehistory of JMM.

Items considered for JMM include:

1. Monographs on mathematics

2. Special lectures on new fields of mathematics

3. Mathematical seminar reports

It is our wish that the publication of the Josai Mathematical Monographs will be fruitful.

February 2001

Hiroyuki Ishibashi

Professor

Dean of the Graduate School of Science Josai University